OpenStudy (anonymous):

Once you have a single variable isolated in one of the two equations in a system, the second step of using substitution to solve the system of equations is to

OpenStudy (dmb1215):

ok so if you were to use this 2x+7=21 first you would say im finding the x then -7 from both sides so 2x=14 and divide it by two and get 7

OpenStudy (dmb1215):

is that what you are meaning

OpenStudy (anonymous):

So would i substitute the expression equal to the solved variable into its original equation?

OpenStudy (dmb1215):

give me a equation that has to do with this plz

OpenStudy (anonymous):

Once you have a single variable isolated in one of the two equations in a system, the second step of using substitution to solve the system of equations is to __________. A. solve for the second variable in the other equation B. solve for the second variable in the same equation C. substitute the expression equal to the solved variable into the other equation D. substitute the expression equal to the solved variable into its original equation

OpenStudy (dmb1215):

@SolomonZelman

OpenStudy (anonymous):

thats the question ^^^^

OpenStudy (solomonzelman):

I think C, because what you are doing is (my own example) y=2x+1 2x+2y=7 2x+2(2x+1)=7

OpenStudy (anonymous):

thank you what about this one?

OpenStudy (anonymous):

In the system –12c + 3d = 6 and –3c + 6d = –9, which expression could you use for substitution?

OpenStudy (dmb1215):

thanks for helping @SolomonZelman

OpenStudy (solomonzelman):

Do yuo have the choices?

OpenStudy (anonymous):

A. d=-1/2c-3/2 B.c=-2d+3 C.c=1/4d-1/2 D. d=4c-2